OpenStudy (theedwardsfamily):

Which statement correctly compares 11/20 and 5/9 ? A. 11/20 < 5/9 B. 11/20 > 5/9 C. 11/20= 5/9

OpenStudy (andrewthecookie):

A i believe

OpenStudy (mertsj):

Change both fractions so the denominators are 180 and then you will see.

OpenStudy (anonymous):

d. 9+10

OpenStudy (anonymous):

@Mertsj why is your profile pic a club penguin membership logo?

OpenStudy (andrewthecookie):

11/20 x 9 = 180 and 5/9 x 20 = 180

OpenStudy (andrewthecookie):

11/180 = 9/180

OpenStudy (andrewthecookie):

So its B

OpenStudy (andrewthecookie):

11/180 > 9/180*

OpenStudy (theedwardsfamily):

OK I perfectly understand thanks @AndrewTheCookie !

OpenStudy (anonymous):

np

OpenStudy (andrewthecookie):

Good :D

OpenStudy (mertsj):

\[\frac{11}{20}\times \frac{9}{9}=\frac{99}{180}\]

OpenStudy (theedwardsfamily):

you too @TheShadyLP !:D

OpenStudy (andrewthecookie):

@Mertsj How do u make the fractions like that?

OpenStudy (anonymous):

ya np @TheEdwardsFamily

OpenStudy (andrewthecookie):

Is it Latex?

OpenStudy (anonymous):

no it is equation @AndrewTheCookie

OpenStudy (andrewthecookie):

oh

OpenStudy (mertsj):

\[\frac{5}{9}\times \frac{20}{20}=\frac{100}{180}\]

OpenStudy (andrewthecookie):

thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

it it my duty to help and serve others of the cookie community

OpenStudy (anonymous):

.........and then eat them..........................

OpenStudy (theedwardsfamily):

@TheShadyLP Lol!!

OpenStudy (mertsj):

So choose A

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

why is your profile pic a club penguin membership logo @Mertsj ???

OpenStudy (mertsj):

I didn't know it was.

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

oh well

OpenStudy (anonymous):

you get a medal for that @Mertsj

OpenStudy (mertsj):

Someone found it for me and said I should use it so I did. A long time ago.

OpenStudy (anonymous):

ok cooooollllliiiiioooo